OpenStudy (anonymous):

What is the slope of the line which passes through (4, 5) and (0, 1)? A. 0 b.1 c. 4 d. undefined

jimthompson5910 (jim_thompson5910):

Subtract the y coordinates. Then subtract the x coordinates (keep the order the same). Then divide the y difference over the x difference.

OpenStudy (anonymous):

That would come out to be 0. Thank you

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

I'm going to subtract the second coordinate from the first y difference : 5 - 1 = 4 x difference: 4 - 0 =4 Slope = (y difference)/(x difference) = 4/4 = 1 So the slope is 1

OpenStudy (anonymous):

Oh okay thanks!!

jimthompson5910 (jim_thompson5910):

yw